### v3

# RankMamba: Benchmarking Mamba’s Document Ranking Performance in the Era of Transformers

Abstract
Transformer structure has achieved great success in multiple applied machine learning communities, such as natural language processing (NLP), computer vision (CV) and information retrieval (IR). Transformer architecture’s core mechanism — attention requires $O(n^{2})$ time complexity in training and $O(n)$ time complexity in inference. Many works have been proposed to improve the attention mechanism’s scalability, such as Flash Attention and Multi-query Attention. A different line of work aims to design new mechanisms to replace attention. Recently, a notable model structure — Mamba, which is based on state space models, has achieved transformer-equivalent performance in multiple sequence modeling tasks. In this work, we examine Mamba’s efficacy through the lens of a classical IR task — document ranking. A reranker model takes a query and a document as input, and predicts a scalar relevance score. This task demands the language model’s ability to comprehend lengthy contextual inputs and to capture the interaction between query and document tokens. We find that (1) Mamba models achieve competitive performance compared to transformer-based models with the same training recipe; (2) but also have a lower training throughput in comparison to efficient transformer implementations such as flash attention. We hope this study can serve as a starting point to explore Mamba models in other classical IR tasks. Our code implementation is made public to facilitate reproducibility. Refer to Xu et al. (2025d) for more comprehensive experiments and results, including passage ranking.
1 Introduction
Transformer structure has been the status quo within NLP and IR community since the introduction of Transformers (Vaswani et al., 2017) and BERT (Devlin et al., 2019). Compared to the classical RNN structure, the transformer structure is better at capturing long-range dependencies and also can be pre-trained on a large scale. However, in inference time, the transformer structure requires $O(n)$ time complexity and $O(n\cdot d)$ space complexity to store KV-cache, where $n$ denotes sequence length and $d$ denotes the size of the hidden state, making it less efficient compared to RNN. In addition, the transformer structure is also hard to extrapolate to context length longer than its pre-training length.
Recently, there has been a growing interest in developing alternative structures for modeling sequence data. For example, RWKV (Peng et al., 2023) combines the efficient parallelizable training of transformers with the efficient inference of RNNs. Another notable structure is state space models (SSM) (Gu and Dao, 2023), which are related to CNN, RNN and the classical state space models. In essence, state space models compress the context into a smaller state, achieving $O(1)$ time complexity and $O(d)$ space complexity in inference time. State space models are efficient in inference but also limited by the amount of information that can be compressed, i.e. the hidden state size. To mitigate this drawback, Gu and Dao (2023) propose a novel selective state space model named Mamba. Mamba   addresses the computation problem of SSM with a selective scan method and corresponding hardware-aware optimization techniques, achieving performance close to transformer-based models of similar sizes while also demonstrating efficiency in training and inference.
Modern search systems typically consist of at least two stages: document retrieval and document ranking. During document retrieval, offline-built indexes are utilized to fetch a preliminary list of candidate documents. Subsequently, the document ranking task refines this initial retrieval list by modeling fine-grained query-document interactions. Document ranking task requires the capability of language models (LM) to understand long context input and to capture the interaction between query and document tokens, at which transformer structure is good by construct, as the attention mechanism allows the query tokens to attend to document tokens. Given the success of Mamba in other sequence modeling tasks, we pose the following research question:
RQ: Can state space models (e.g., Mamba) achieve comparable performance to transformer-based language models in the document ranking task?
To comprehensively assess the performance of Mamba, we conduct a rigorous benchmarking study against transformer-based models characterized by varying structures, sizes, pre-training objectives, and attention patterns. Correspondingly, we train neural reranker models following established training methodologies outlined in prior literature (Boytsov et al., 2022; Gao et al., 2021; Ma et al., 2023) This comparison allows us to evaluate the capabilities of different LMs in terms of performance and training efficiency. We find:
  • Encoder-only transformer-based LMs continue to demonstrate competitive performance in the document ranking task. For instance, roberta-large achieves 0.4434 MRR on MSMARCO Dev set, outperforming encoder-decoder LMs and decoder-only LMs of comparable size.
  • Mamba models can achieve competitive performance, often matching or even surpassing transformer-based LMs of similar sizes. However, they also have lower training throughput compared to more efficient transformer implementations.
We first discuss the background of Mamba models in Section 2. We introduce our experiment setup in Section 3 and report the experimental results in Section 4. We review related works in Section 5, conclude this work and discuss limitations in Section 6.
2 Background on Selective State Space Models
2.1 Structured State Space Sequence Models (S4)
Prior works on state space models (Gu et al., 2021b; Gupta et al., 2022; Gu et al., 2022; Smith et al., 2022) are linear time invariance (LTI). We use Structured State Space Sequence Models (S4) Gu et al. (2021a) as a representative example to illustrate the idea behind vanilla state space models.
State space models map a 1-dimensional function or sequence $x(t)\in\mathbb{R}$ to $y(t)\in\mathbb{R}$ through a latent state $h(t)\in\mathbb{R}^{N}$, where $t$ denotes a timestep. They are parameterized by $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ and define a continuous sequence-to-sequence transformation as:
[…]
The above transformation can be discretized as:
$h_{t}$ $=\overline{\textbf{\emph{A}}}h_{t-1}+\overline{\textbf{\emph{B}}}x_{t}$ $y_{t}$ $=\textbf{\emph{C}}h_{t}$ (2)
The discretization of A and $\overline{\textbf{\emph{B}}}$ is defined by the discretization rule, for example:
$\overline{\textbf{\emph{A}}}$ $=\exp(\Delta\textbf{\emph{A}})$ $\overline{\textbf{\emph{B}}}$ $=(\Delta\textbf{\emph{A}})^{-1}(\exp(\Delta\textbf{\emph{A}})-I)\cdot\Delta\textbf{\emph{B}}$ (3)
If we expand Equation 2 with the whole sequence $x=(x_{1},x_{2},\ldots,x_{n})$, we can arrive at:
$\overline{\textbf{\emph{K}}}$ $=(\textbf{\emph{C}}\overline{\textbf{\emph{B}}},C\overline{\textbf{\emph{A}}\textbf{\emph{B}}},\ldots,C\overline{\textbf{\emph{A}}}^{n-1}\overline{\textbf{\emph{B}}})$ $y$ $=x*\overline{\textbf{\emph{K}}}$ (4)
Readers might have noticed that Equation 2 looks like RNN while Equation 4 looks like CNN, where $\overline{\textbf{\emph{K}}}$ is a large convolution kernel over the whole input sequence $x$. The parameterization of $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ is independent of input sequence $x$ and is fixed during all time steps, a property referred to as linear time invariance (LTI). Structured state space models (S4) impose structure on the A matrix to be computed efficiently. [^1] An important note is that A matrix also requires a specific initialization method: HiPPO Gu et al. (2020) to memorize the context, the details of which will not be covered in this work. Existing works (Gu et al., 2021a; Gupta et al., 2022; Smith et al., 2022) mostly deploy diagonal matrix, thus $\textbf{\emph{A}}\in\mathbb{R}^{N\times N}$, $\textbf{\emph{B}}\in\mathbb{R}^{N\times 1}$ and $\textbf{\emph{C}}\in\mathbb{R}^{1\times N}$ matrices are all represented by $N$ parameters.
Equation 1-4 can be generalized to $D$-channel features, i.e. $x_{t},y_{t}\in\mathbb{R}^{D}$, a concrete example might be a word embedding of dimension $D$. In this case, computation of $\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}}$ is applied to each channel independently.
2.2 Selective Scan Structured State Space Sequence Models (S6)
State space models compress context into a smaller hidden state $h_{t}\in\mathbb{R}^{N}$, which limits the model’s effectiveness. Gu and Dao (2023) propose to modify a subset of parameters $(\Delta,\textbf{\emph{B}},\textbf{\emph{C}})$ as input-dependent. This modification changes the model from time-invariant to time-varying, and poses challenges to the model’s computational efficiency, as the model now cannot be trained in CNN mode. Specifically, denote the context length as $L$ and batch size as $B$, training in RNN mode yields a total FLOPs of $O(BLDN)$, which cannot be trivially parallelized as each step depends on the output of the previous step.
To tackle the computation problem, Gu and Dao (2023) propose a hardware-aware optimization algorithm named Selective Scan. It mainly comprises the following critical improvements:
  • Instead of computing $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}})$ of size $(B,L,D,N)$ in GPU HBM (High-Bandwidth Memory), now SSM parameters $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ are loaded in GPU fast SRAM (Static Random Access Memory) to speed up the computing, and output of size $(B,L,D)$ are written back to HBM. This optimization technique is similar to Flash Attention Dao (2024).
  • To avoid the sequential recurrence and to perform the computation in parallel, a work-efficient parallel scan algorithm Blelloch ; Martin and Cundy (2018) is used to amortize the computation to multiple threads within GPU.
  • Instead of saving the intermediate states, they are recomputed in the backward pass when the inputs are loaded from HBM to SRAM, which vastly reduces the memory requirements.
Gu and Dao (2023) parameterize the corresponding selective state space model into a Mamba block, in which the selective scan structured state space sequence (S6) model only serves as a replacement for Attention component, while the majority of computation is handled by FFN layers and corresponding activation layers. Combining the aforementioned optimization techniques, the parameterized Mamba block can be computed efficiently and achieve 5× higher throughput on long sequence lengths compared to Transformers of similar sizes, as reported in Gu and Dao (2023).
3 Methodology and Experiment Setup
We have briefly discussed state space models and the design choices of Mamba model. In this section, we focus on evaluating Mamba’s efficacy in the document ranking task. To this end, we aim to address the following research questions:
  • RQ1: Between Mamba and transformer-based models of similar sizes, can Mamba perform comparably in document ranking task?
  • RQ2: Can Mamba achieve higher training throughput compared to state-of-the-art efficient attention implementations?
Given the diverse array of existing transformer-based language models, we mainly focus on three distinct types: encoder-only models, decoder-only models, and encoder-decoder models. We elaborate on our design choices in this section.
3.1 Task Formulation
We focus on the reranking task, where the input to the reranker model is a search query and a candidate document from a first-stage retriever, and the model returns a scalar score indicating the relevance between the query and the document.
Denote the reranker model as $f$, input query as $q\in\mathcal{Q}$, input document as $d\in\mathcal{D}$ and predicted relevance score as $s$,
$f(q,d)\rightarrow s\in\mathbb{R}$ (5)
Depending on the backbone language models we use, the exact ways to format $(q,d)$ are different. For models pre-trained with bi-directional attention, we prepend a [CLS]token to the input. On top of the backbone model, a fully connected layer (FFN) is randomly initialized, which takes the representation of the [CLS]token as input and predicts the relevance score $s$. The input format is
[…]
where [SEP]and [EOS]are special tokens indicating separation and end of sequence, respectively.
Encoder-decoder models, i.e. T5   family models can be used for feature extraction such as encoder-only models, and can be used for text generation such as decoder-only models. As the encoder part is pre-trained with bidirectional attention, we use the same input format as encoder-only models. Then the FFN layer takes the representation of the first decoded token and predicts the relevance score. This setup is the same as prior works on adapting T5 for the document ranking task (Zhuang et al., 2023; Ni et al., 2022).
For decoder-only models pre-trained with unidirectional attention (often referred to as autoregressive LM), we use the following input format
[…]
The intuition for putting query $q$ at the later part of the input is to allow the query tokens to attend to prior document tokens to better capture relevance. On top of the backbone autoregressive LM, the randomly initialized FFN layer takes the representation of the [EOS] token as input and predicts the relevance score $s$.
For Mamba models, as they are pre-trained with unidirectional information, we use the same input format as causal LMs.
Table 1: We list the compared backbone language models in terms of size, information, pre-training objective as well as model structure. NSP refers to the next sentence prediction objective. $\mathbf{a}$ Note that T5   family models are additionally multitask fine-tuned with GLUE tasks Wang et al. (2018).
Model Size Information Pre-training Objective Model Structure Positional Encoding
Encoder-only Models
bert-base-uncased 110M Bidirectional MLM & NSP Attention Learned
bert-large-uncased 330M Bidirectional MLM & NSP Attention Learned
roberta-base 125M Bidirectional MLM Attention Learned
roberta-large 355M Bidirectional MLM Attention Learned
Decoder-only Models
facebook/opt-125m 125M Unidirectional CLM Attention Learned
facebook/opt-350m 350M Unidirectional CLM Attention Learned
EleutherAI/pythia-160m 120M Unidirectional CLM Attention Rotary
EleutherAI/pythia-410m 340M Unidirectional CLM Attention Rotary
EleutherAI/pythia-1b 860M Unidirectional CLM Attention Rotary
state-spaces/mamba-130m 130M Unidirectional CLM Mamba None
state-spaces/mamba-370m 370M Unidirectional CLM Mamba None
state-spaces/mamba-790m 790M Unidirectional CLM Mamba None
Encoder-Decoder Models
T5-base 220M Bidirectional Denoisinga Attention Relative
T5-large 770M Bidirectional Denoising Attention Relative
3.2 Training Objective
Denote the relevant document to query $q_{i}$ as $d_{i}^{+}$ and sampled negative documents as $d_{i}^{-}\in\mathcal{D}_{i}^{-}$, the reranker is trained by optimizing a InfoNCE loss
[…]
In practice, each mini-batch consists of multiple queries and corresponding documents, and loss is averaged over each query in the mini-batch before the optimizer takes an optimization step.
3.3 Backbone Language Models
Table 2: We list the training details and choice of hyperparameters. LR denotes learning rate, AMP denotes PyTorch Automatic Mixed Precision Training. Note that LoRA training requires a larger learning rate.
Model Size LR AMP Flash Attention LoRA
Encoder-only Models
bert-base-uncased 110M 2e-5 FP16
bert-large-uncased 330M 1e-5 FP16
roberta-base 125M 2e-5 FP16
roberta-large 355M 1e-5 FP16
Decoder-only Models
facebook/opt-125m 125M 2e-5 BF16
facebook/opt-350m 350M 1e-5 BF16
EleutherAI/pythia-160m 120M 2e-5 BF16
EleutherAI/pythia-410m 340M 1e-5 BF16
EleutherAI/pythia-1b 860M 1e-4 BF16
state-spaces/mamba-130m-hf 130M 2e-5 BF16
state-spaces/mamba-370m-hf 370M 1e-5 BF16
state-spaces/mamba-790m-hf 790M 1e-4 BF16
Encoder-Decoder Models
T5-base 220M 2e-5 BF16
T5-large 770M 1e-4 BF16
Table 3: We report official evaluation metrics for MSMARCO Dev, TREC DL19 and DL20. The highest number in each section is highlighted.
MS MARCO Dev TREC DL19 TREC DL20
Model Size MRR NDCG@10
BM25(k1=4.46, b=0.82) - 0.2767 0.5233 0.5061
bert-base-uncased 110M 0.4126 0.6575 0.6147
roberta-base 125M 0.4006 0.6591 0.6140
facebook/opt-125m 125M 0.3878 0.6377 0.6176
EleutherAI/pythia-160m 120M 0.3727 0.6490 0.6362
state-spaces/mamba-130m 130M 0.4089 0.6652 0.6443
T5-base 220M 0.3670 0.6383 0.5932
bert-large-uncased 330M 0.4006 0.6591 0.6140
roberta-large 355M 0.4334 0.6676 0.6423
facebook/opt-350m 350M 0.3569 0.6431 0.6046
EleutherAI/pythia-410m 340M 0.3750 0.6497 0.6169
state-spaces/mamba-370m 370M 0.4250 0.6775 0.6394
T5-large 770M 0.3881 0.6473 0.6146
EleutherAI/pythia-1b 860M 0.3921 0.6588 0.6386
state-spaces/mamba-790m 790M 0.4201 0.6729 0.6489
We benchmark the performance of different models on three sizes: models with 110M parameters (e.g. bert-base-uncased); models with 330M parameters (e.g. bert-large-uncased) and models with >> 700M parameters (e.g. T5-large and state-spaces/mamba-790m).
For encoder-only models, we choose BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019); for decoder-only models, we include Pythia (Biderman et al., 2023), OPT (Zhang et al., 2022) and Mamba (Gu and Dao, 2023), and we opt for T5 (Raffel et al., 2020) as the representative encoder-decoder models. The compared models are from different model families, varying in terms of pre-training objectives (e.g. causal language modeling, masked language modeling, denoising, next sentence prediction), information direction (bidirectional vs unidirectional), model structure (Attention vs Mamba) and Positional Encoding (no positional encoding, learned positional encoding, relative positional encoding and rotary positional encoding (Su et al., 2024)). An important note is that we do not include instruction-finetuned language models such as Flan-T5 models (Chung et al., 2022) for a fair comparison to other language model families. For a review of these models, refer to Table 1.
3.4 Datasets
We use MS MARCO document ranking dataset for training. The dataset consists of 3.2M documents sampled from search results of Bing search engine. We use the training queries and qrels from the official training set.
Following prior practices (Gao et al., 2021; Boytsov et al., 2022; Ma et al., 2023), for each query and relevant document from the training set, we sample negative documents from the top-100 results returned by a sparse BM25 retriever implemented by Pyserini [^2] https://github.com/castorini/pyserini (Lin et al., 2021). In practice, we fix the random seed and sample 72 negative documents for each qrel in the official training set, leading to in total 734,008 training samples, each consisting of 1 relevant document and 7 hard negative documents. Notably, we use the setup recommended by Gao et al. (2021) to sample hard negatives from the top-100 results returned by the first-stage retriever (BM25 in our case). The sample training set is used throughout all experiments to ensure fair comparison and reproducibility.
We use the official document ranking Dev set for evaluation. We also use TREC DL19 (Craswell et al., 2020) and DL20 (Craswell et al., 2021). The official Dev set consists of 5,193 queries, with 1 relevant judgment each, while DL19 and DL20 consist of 43 and 45 queries respectively with denser relevance judgments.
We use the reranking setup in evaluation as prior practices (Boytsov et al., 2022; Gao et al., 2021; Ma et al., 2023) where a first-stage retriever BM25 retrieves top-100 documents and our trained reranker models perform the reranking. We report Mean Reciprocal Rank@100 for Dev set and NDCG@10 for DL19 and DL20 to be consistent with the official evaluation. Deploying better first-stage retrievers can lead to better reranking performance but is beyond the scope of this study.
3.5 Training Details and Hyperparameters
We truncate the input to 512 tokens to fit in the context window of BERT and RoBERTa and T5 models (also referred to as FirstP by Dai and Callan (2019)). For autoregressive LMs (OPT, Pythia and Mamba), although they have a longer context length, we also use 512 tokens to keep a fair comparison. We use the official implementation of Mamba models [^3] https://github.com/state-spaces/mamba and weights of other models are downloaded from Huggingface. [^4] https://huggingface.co/models
For all experiments, we optimize the reranker model with AdamW optimizer (Loshchilov and Hutter, 2019), with betas=(0.9, 0.999), eps=1e-08, weight_decay=0.01. In our preliminary experiments, we noticed models larger than 300M can suffer from overfitting, therefore we use a smaller learning rate for larger models. We use a linear warmup scheduling where we gradually increase the learning rate in the first 1000 training steps and decrease to zero until the end of training. We use PyTorch native mixed precision training and Flash Attention 2 Dao (2024) for higher throughput. For models larger than 700M, i.e. T5-large, EleutherAI/pythia-1b and state-spaces/mamba-790m, we use LoRA (Hu et al., 2021) to reduce memory consumption and accelerate training.
We train the reranker models for 1 single epoch, as recommended by Ma et al. (2023). Through the course of training, for each reranker model, we save 10 intermediate checkpoints to monitor the training trajectory and report the performance of the final checkpoint. Refer to Table 2 for a detailed listing of hyperparameters and training details.
4 Result and Analysis
Document Ranking Performance. We report the ranking performance in Table 3. We notice that encoder-only language models such as BERT and RoBERTa can achieve strong performance. For example, within models 110M and models 330M parameters, bert-base-uncased and roberta-large achieve the highest MRR on MSMARCO Dev dataset.
For decoder-only LMs such as OPT and Pythia, we notice an interesting discrepancy between performance on datasets with in-depth relevance judgments (DL19 and DL20) and MSMARCO Dev which has shallow judgments (1 relevant document per query). A similar trend also applies to T5 family models. We conjecture this discrepancy might be due to different annotation schemes adopted by MSMARCO Dev and TREC DL, and will leave this to potential future work.
With the current training recipe, Mamba models achieve competitive performance compared to transformer-based models of similar scales. For example, among models 110M parameters, state-spaces/mamba-130m achieves the highest NDCG@10 on TREC DL19 and DL20 datasets. state-spaces/mamba-370m also achieves highest NDCG@10 on DL19, and second highest NDCG@10 on DL20 among 5 models we compare with scale of 330M parameters. Mamba’s strong performance is consistent when trained with LoRA. For example, state-spaces/mamba-790m compared to T5-large and EleutherAI/pythia-1b, approaching full parameter fine-tuning of state-spaces/mamba-370m.
Benchmarking Training Throughput for Document Ranking.
We benchmark the training throughput of different models on a server with one A6000 GPU and report results in Figure 1. Overall, models trained with flash attention implementation (OPT, Pythia) achieve the highest training throughput and lowest space complexity. We notice that although Mamba can achieve $O(n)$ training complexity in theory, the training throughput with its current implementation is still lower than efficient transformer implementations such as flash attention. This finding suggests the deficiency of Mamba’s current implementation and leaves us with potential future work for improvement.
Figure 1: We show the training throughput of models 330M parameters and models >> 700M parameters. Models >> 700M parameters are trained with LoRA, with rank=32. We notice Mamba models have lower throughput and higher GPU memory consumption compared to efficient transformer implementations such as flash attention.
5 Related Works
Transformer-based Language Models
These language models utilize the attention mechanism (Vaswani et al., 2017) to produce contextualized representations of tokens and demonstrate superior performance and training efficiency compared to traditional recurrent neural network models (Hochreiter and Schmidhuber, 1997; Peters et al., 2018). Such LMs can be used as feature extractors for downstream tasks or be used for natural language generation (NLG) with a specific language modeling head. Prior works (Xu and Jiang, 2024; Raffel et al., 2020; Wang et al., 2022; Le Scao et al., 2022) have shown that between models of similar sizes, different pre-training objectives can lead to strengths and weaknesses in terms of natural language understanding (NLU) and natural language generation capabilities.
Language Models in IR
Neural network-based language modeling techniques have gained widespread adoption within the information retrieval (IR) community, including web search snippet generation (Bast and Celikik, 2014; Xu and Cohen, 2023; Xu, 2023), modeling users and products in recommender systems (Xu et al., 2021; Zeng et al., 2021), generating search results explanation (Yu et al., 2022; Xu et al., 2023) and conversational search (Wang et al., 2023b, a; Aliannejadi et al., 2019).
Document Ranking
The document ranking task aims to capture the relevance between query and document. Traditional bag-of-words methods such as BM25 (Robertson et al., 1999) aim to capture the lexical overlapping. Neural models such as word2vec (Mikolov et al., 2013) are learned from static word co-occurrence patterns from a pre-training corpus and can capture the semantic meanings of the words. Transformer-based language models can capture contextualized representations and have been widely applied to modeling query-document relevance. Dai and Callan (2019) propose to model long documents with BERT (Devlin et al., 2019). To model long documents with BERT’s limited 512 context window, they propose several model variants such as FirstP, MaxP and SumP. Followup works (Li et al., 2023; Hofstätter et al., 2021; Boytsov et al., 2022; Ma et al., 2023) achieve performance improvements by leveraging techniques such as mining hard negatives, chunking long documents and/or sliding windows, and replacing BERT with better, long-context pre-trained language models such as Longformer (Beltagy et al., 2020) and Llama (Touvron et al., 2023a, b). For a more comprehensive literature review, refer to Lin et al. (2022) and Xu et al. (2025c)
6 Conclusion and Limitations
This study aims to benchmark Mamba models’ performance in the document ranking task. We find that Mamba models can achieve competitive performance compared to transformer-based language models, and the effectiveness is consistent with trained with LoRA. However, Mamba models have the deficiency of lower training throughput compared to efficient attention implementations, limiting their potential for efficient training and deployment. We only explored the contrastive learning training strategy. As recent works demonstrate, knowledge distillation is especially effective for training lightweight rerankers when distilling from a more capable reranker teacher (Baldelli et al., 2024; Schlatt et al., 2025; Xu et al., 2025b). Due to the limited bandwidth, we only study reranker models << 1B parameters based on the BM25 retriever. Our future work should include extending to rerankers of large scales, different document retrievers.